OpenStudy (deadshot):

How do I solve 4^x = 12?

OpenStudy (deadshot):

\[4^x = 12\]

OpenStudy (anonymous):

Its asking you 4^(to the power of what)= 12

OpenStudy (anonymous):

Here's a example: 4^x = 32 (2^2)^(x) = 2^5 2 ^(2x) = 2^5 2x = 5 x = 2.5

OpenStudy (anonymous):

Hope this helps! :)

OpenStudy (anonymous):

@zerosniper123 it seems this method is difficult to apply in this problem

OpenStudy (anonymous):

Im sorry I never completed the rest of this, I'm busy with helping others as well. I was just giving him a example, I had not yet finished detailing since that this problem doesn't have a perfect whole number root that he will have to use another method.

OpenStudy (anonymous):

u can aply log on both sides and and clear like: \[4^{x}=12\] \[\log(4^{x})=\log(12)\] now log propierties = xlog(4)=log(12) then x=log4/log12

OpenStudy (anonymous):

@zerosniper123 its okay buddy @Kewwszz \[x = \frac{ \log12 }{ \log4 }\]